OpenStudy (anonymous):

I NEED HELP PLEASE IN MATH Part 1 - A family of 5 comes to the amusement park. There are two adults and three children (under the age of 12). They qualify for a special family rate of $174.45. The clerk tells them that a child's ticket always costs $11 less than an adult's ticket. How much is each adult ticket and how much is each child's ticket? 1.Define your variables. What is a variable you can use to represent the adult cost? A variable to represent adult cost in the family plan: __________.

OpenStudy (anonymous):

2a + 3c = 174.45 c = a -11 used c for child ticket and a for adult ticket. x and y can also be used instead. substitute c = a - 11 in first row and you'll get an answer

OpenStudy (anonymous):

Let the rate of an adult ticket=x then rate of a child ticket=x-11 Also 2 x+3(x-11)=174.45 or 2 x+3x -33=174.45 or 5x=174.45+33=207.45 or x=207.45/5=41.49$ therefore rate of adult ticket=41.49$ rate of chid's ticket=41.49-11=30.49$